# Taming generative video models for zero-shot optical flow extraction

### v1

@@ 4.2 Study 2: Diffusion models’ global latent code lacks sufficient fine-grained controllability
Method: Applying the method above (Section 2, Figure 2) to Stable Video Diffusion [5] is challenging as generation operates on full-frame latents rather than pixels, so these models do not natively support partial frames as input. However, we need to provide information from the second frame to anchor the generation, otherwise the model will hallucinate arbitrary future frames unsuitable for flow estimation. Below, we describe how we adapt the flow extraction procedure to the latent diffusion setting, while keeping the generations anchored to the observed video. Step 1. Frame selection and perturbation. As SVD is trained with a fixed framerate and more than two frames, we first retrieve an interpolated sequence of intermediate frames bridging the query ($F_{1}$) and target ($F_{N}$) frames =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{N}\}$. We add a dot perturbation to $F_{1}$, obtaining $\tilde{F}_{1}$. Step 2. Latent inversion and paired generation. To anchor the sampling trajectory to the original video, we apply a latent inversion technique introduced in VideoShop [10]. Inversion addresses the drift in generation by finding an initial noise vector that reconstructs the input video when fed into the diffusion process. Specifically, we first obtain the latents for every frame =$\boldsymbol{\ell}=\{\ell_{1},\ell_{2},\dots,\ell_{N}\}$ and then partially apply the inversion process, adding noise to the latents up to a fraction of the total noising steps. This is done using SVD’s UNet and an inverted scheduler, resulting in noisy latents ~={~$\tilde{\boldsymbol{\ell}}=\{\tilde{\ell}_{1},\dots,\tilde{\ell}_{N}\}$. From these latents we perform two denoising passes with identical hyperparameters, but with different conditioning frames, $F_{1}$ and the perturbed $\tilde{F}_{1}$, obtaining two denoised latents $\boldsymbol{\ell}^{\text{pred}}$ and ~$\boldsymbol{\tilde{\ell}}^{\text{pred}}$, which can be decoded to produce the clean $\mathcal{F}^{\text{pred}}$and perturbed reconstruction ~$\tilde{\mathcal{F}}^{\text{pred}}$. Step 3. Estimate optical flow (same as in Section 2). We compute the RGB difference between the clean and perturbed generations to localize the perturbation. We compare two generations, rather than the perturbed generation to the original input, in order to suppress noise from imperfect inversion and VAE artifacts, improving the accuracy and robustness of flow predictions even under lossy latent compression.
@@
Attempting to address the precision issues arising from a coarse global latent, the recent DiffTrack [28] uses a one-to-one frame-to-latent mapping to avoid temporal compression. It uses query-key matching in the 3D attention blocks across frames to perform point tracking. Nevertheless, DiffTrack performs worse than specialized flow models (Table 3). Next, we explore autoregressive models which allow us to provide actual patches of the second frame as more fine-grained conditioning.
@@ 4.3 Study 3: Raster-order autoregressive models struggle with partial-frame conditioning
Method: Cosmos [29] includes both a diffusion-based and an autoregressive foundation model. We evaluate flow extractions from the autoregressive model which allows us to provide the actual frame patches directly as conditioning. We use the Cosmos autoregressive model with 4B parameters, which comes with a 7B diffusion decoder for generating images. As the Cosmos autoregressive model does not have pointer tokens, autoregressive rollout predictions for image tokens cannot be performed in random-access order. Instead, they can only be made in raster order, i.e., from top left to bottom right of the image. Nevertheless, we make a best-effort attempt to extract flow by trying multiple settings.
@@ 4.4 Study 4: Key model properties for flow extraction are present in LRAS
Method: From Sections 4.1, 4.2, and 4.3, we discover that flow extraction requires generative models to have three key properties (Table 1): (1) distributional prediction of future frames (avoiding blurry or noisy outputs; which CWM lacks); (2) local tokenizer that treats each spatio-temporal patch independently (which SVD and Cosmos lack); and (3) random-access decoding order that allows the model to condition on any subset of the second frame patches (which SVD and Cosmos lack). We identify Local Random Access Sequence (LRAS), a recent family of autoregressive visual foundation world models that has these properties and can be zero-shot prompted to perform many tasks, such as 3D object manipulation, novel view synthesis, and depth estimation. We use its autoregressive generative video world model with 7 billion (7B) parameters.
@@ 5 KL-tracing bypasses sampling randomness
Figure 4: Our method, KL-tracing using LRAS extracts better flow than other generative video models. (A) Deterministic models, such as CWM [3], often produce blurry predictions as they model a single, average future state. (B) Stable Video Diffusion lacks fine-grained controllability due to its coarse global latent code. Its clean and perturbed predictions differ in locations where the perturbation is not supposed to be carried to. (C) The Cosmos autoregressive world model lacks fine-grained controllability as it does not utilize pointers to denote the position of each token, making it challenging to prompt for flow extraction. (D) The LRAS model is highly controllable and has minimal differences between the clean and perturbed predictions. We use KL-tracing to compute the difference in logit instead of RGB space, obtaining sharp flow extractions.
@@
Figure 6: Large world models such as LRAS capture aspects of real-world dynamics that are challenging for specialized flow models relying on visual similarity or photometric loss. Specialized optical flow baselines, both supervised and self-supervised, struggle on complex, real-world scenes that are not fully captured by their training heuristics (Section 6). A deeper understanding of dynamics beyond feature/photometric similarity allows large world models such as LRAS to resolve motion in the presence of homogeneous objects, motion blur, and partial occlusion, among others.
6 Large video world models capture aspects of real-world dynamics that are challenging for specialized flow models
Method: We compare KL-tracing with supervised flow methods such as RAFT [37] and SEA-RAFT [39], as well as unsupervised methods such as Doduo [21] and SMURF [36] (Table 3, Figure 6).
Findings: Powerful video models capture challenging aspects of real-world dynamics. For example, a globe undergoes in-place rotation with the query point on a textureless surface (Figure 1), or physical motion examples such as Newton’s cradle (Figure 6). These real-world dynamics are challenging for specialized flow models which are small and not trained on diverse, large-scale video datasets. Video world models use generic training objectives, so they do not suffer from failure modes introduced by training heuristics used by specialized flow models. Supervised methods are trained on synthetic datasets as labeling real-world videos is expensive, and hence face a sim-to-real gap where they do not observe challenging real-world dynamics. Self-supervised methods such as Doduo [21] and SMURF [36] use photometric loss for training, which results in challenges in predicting flow for frame pairs with significant differences in light intensity, and also enforce strong global consistency and smoothness constraints, which introduce failure modes for complex non-uniform dynamics that contain various magnitudes of flow across the same frame pair.
CWM [3] SVD [5] Cosmos [29] LRAS [24]
@@
Stable Video Diffusion [5] 74.7990
Cosmos (top 10% raster) (5MM, 2STD, 512$\times$512) [29] 35.4338
Cosmos (overwrite 10% during rollout) (5MM, 2STD, 512$\times$512) [29] 37.7552
Cosmos (provide full second frame) (5MM, 2STD, 512$\times$512) [29] 66.5521
Table 2: KL-tracing with LRAS beats other video models. MM = multi-mask (average over multiple random masks), MS = multi-scale (zoom into image at multiple scales).
@@
AJ $\uparrow$ AD $\downarrow$ $<\!\delta^{x}_{\text{avg}}\!$uparrow OA $\uparrow$ AJ $\uparrow$ AD $\downarrow$ $<\!\delta^{x}_{\text{avg}}\!$uparrow OA $\uparrow$
@@
S RAFT [37] 41.77 25.33 54.37 66.40 71.93 5.60 82.15 88.54
SEA-RAFT [39] 43.41 20.18 58.69 66.34 75.06 6.54 84.63 89.50
W Doduo [21] 23.34 13.41 48.50 47.91 54.98 5.31 72.20 73.56
U SMURF [36] 30.64 27.28 44.18 59.15 65.81 6.81 80.57 87.91
DiffTrack [28] - - 46.90 - - - - -
CWM [3, 38] 15.00 23.53 26.30 76.63 28.77 11.64 41.63 84.93
@@
Table 3: TAP-Vid First: quantitative results on DAVIS and Kubric. Tracking starts when a point first appears and continues to the video end, thus involving large frame gaps. LRAS with KL-tracing outperforms two-frame baselines. “S" = supervised, “W" = weakly supervised, “U" = unsupervised.
@@ 7 Discussion
Taming generative world models with fine-grained controllability. A key finding of our work is that zero-shot prompting many mainstream generative video models for visual property extraction can be challenging due to their lack of fine-grained controllability. However, a series of design features can help tame these models: (1) predicting a distribution of future states rather than a deterministic average, (2) encoding each spatio-temporal patch independently, and (3) having a mechanism to decode individual parts of a frame in any given order. Currently, the LRAS [24] framework was the only one we identified that contained all of these properties, but they could be adopted by other generative models as well, increasing the ease with which they can be visually prompted.
Future work. (1) A natural next step is extending this test-time inference procedure to extract other visual intermediates such as object segments, material properties or motion affordances. By moving away from fine-tuning our world models on labeled datasets to a paradigm of zero-shot prompting, the vision community can experience a shift akin to the one seen with large language models, moving from rigid representations constructed by fine-tuning toward dynamic, task-specific and adaptable ones extracted on the fly. (2) Extracting flow from large video models is more computationally demanding than a dense flow estimator such as RAFT [37]. However, our model can act as an oracle for high-quality labels. Our next step is distilling LRAS with KL-tracing into a faster architecture.
Overall, our results indicate that prompting controllable, self-supervised world models is a scalable and effective alternative to supervised or photometric-loss approaches for high-quality optical flow.
8 Acknowledgments
This work was supported by the following awards: Simons Foundation grant SFI-AN-NC-GB-Culmination-00002986-05, National Science Foundation CAREER grant 1844724, National Science Foundation Grant NCS-FR 2123963, Office of Naval Research grant N00014-20-1-2589, ONR MURI N00014-21-1-2801, ONR MURI N00014-24-1-2748, and ONR MURI N00014-22-1-2740. We also thank the Stanford HAI, Stanford Data Sciences, the Marlowe team, and the Google TPU Research Cloud team for their computing support.
@@ Appendix A Additional details on methods
CWM [3] White Gaussian Frame 1, Partial Frame 2 RGB difference
SVD [5] White Gaussian Noised latents of Frame 1 and 2 RGB difference
Cosmos [29] White Gaussian Frame 1, Partial Frame 2 RGB difference
LRAS, RGB [24] White Gaussian Frame 1, Partial Frame 2 RGB difference
@@ Appendix B Additional results
Figure 7: All evaluation settings for Cosmos [29]) result in poor flow extractions. See Section 4.3 for a more detailed explanation of each result.

### v2

@@ 4.2 Study 2: Diffusion models’ global latent code lacks sufficient fine-grained controllability
Method: Applying the method above (Section 2, Figure 2) to Stable Video Diffusion [blattmann_stable_2023] is challenging as generation operates on full-frame latents rather than pixels, so these models do not natively support partial frames as input. However, we need to provide information from the second frame to anchor the generation, otherwise the model will hallucinate arbitrary future frames unsuitable for flow estimation. Below, we describe how we adapt the flow extraction procedure to the latent diffusion setting, while keeping the generations anchored to the observed video. Step 1. Frame selection and perturbation. As SVD is trained with a fixed framerate and more than two frames, we first retrieve an interpolated sequence of intermediate frames bridging the query ($F_{1}$) and target ($F_{N}$) frames $\mathcal{F}=\{F_{1},F_{2},\dots,F_{N}\}$. We add a dot perturbation to $F_{1}$, obtaining $\tilde{F}_{1}$. Step 2. Latent inversion and paired generation. To anchor the sampling trajectory to the original video, we apply a latent inversion technique introduced in VideoShop [fan2024videoshop]. Inversion addresses the drift in generation by finding an initial noise vector that reconstructs the input video when fed into the diffusion process. Specifically, we first obtain the latents for every frame $\boldsymbol{\ell}=\{\ell_{1},\ell_{2},\dots,\ell_{N}\}$ and then partially apply the inversion process, adding noise to the latents up to a fraction of the total noising steps. This is done using SVD’s UNet and an inverted scheduler, resulting in noisy latents $\tilde{\boldsymbol{\ell}}=\{\tilde{\ell}_{1},\dots,\tilde{\ell}_{N}\}$. From these latents we perform two denoising passes with identical hyperparameters, but with different conditioning frames, $F_{1}$ and the perturbed $\tilde{F}_{1}$, obtaining two denoised latents $\boldsymbol{\ell}^{\text{pred}}$ and $\boldsymbol{\tilde{\ell}}^{\text{pred}}$, which can be decoded to produce the clean $\mathcal{F}^{\text{pred}}$and perturbed reconstruction $\tilde{\mathcal{F}}^{\text{pred}}$. Step 3. Estimate optical flow (same as in Section 2). We compute the RGB difference between the clean and perturbed generations to localize the perturbation. We compare two generations, rather than the perturbed generation to the original input, in order to suppress noise from imperfect inversion and VAE artifacts, improving the accuracy and robustness of flow predictions even under lossy latent compression.
@@
Attempting to address the precision issues arising from a coarse global latent, the recent DiffTrack [nam2025emergenttemporalcorrespondencesvideo] uses a one-to-one frame-to-latent mapping to avoid temporal compression. It uses query-key matching in the 3D attention blocks across frames to perform point tracking. Though this method demonstrates that strong temporal correspondence can be found in diffusion model representations, with significant improvement over existing diffusion-based methods such as DIFT [tang2023emergentcorrespondenceimagediffusion], it still lags behind specialized optical flow models, thus failing to close the gap between general-purpose video models and task-specific baselines (Table 3). Next, we explore autoregressive models which allow us to provide actual patches of the second frame as more fine-grained conditioning.
@@ 4.3 Study 3: Raster-order autoregressive models struggle with partial-frame conditioning
Method: Cosmos [nvidia_cosmos_2025] includes both a diffusion-based and an autoregressive foundation model. We evaluate flow extractions from the autoregressive model which allows us to provide the actual frame patches directly as conditioning. We use the Cosmos autoregressive model with 4B parameters, which comes with a 7B diffusion decoder for generating images. As the Cosmos autoregressive model does not have pointer tokens, autoregressive rollout predictions for image tokens cannot be performed in random-access order. Instead, they can only be made in raster order, i.e., from top left to bottom right of the image. Nevertheless, we make a best-effort attempt to extract flow by trying multiple settings.
@@ 4.4 Study 4: Key model properties for flow extraction are present in LRAS
Method: From Sections 4.1, 4.2, and 4.3, we discover that flow extraction can be improved by generative models having three key properties (Table 1): (1) distributional prediction of future frames (avoiding blurry or noisy outputs, which CWM lacks); (2) local tokenizer that treats each spatio-temporal patch independently (which SVD and Cosmos lack); and (3) random-access decoding order that allows the model to condition on any subset of the second frame patches (which SVD and Cosmos lack). We identify Local Random Access Sequence (LRAS), a recent family of autoregressive visual foundation world models that has these properties and can be zero-shot prompted to perform many tasks, such as 3D object manipulation, novel view synthesis, and depth estimation. We use its autoregressive generative video world model with 7 billion (7B) parameters.
@@ 5 KL-tracing bypasses sampling randomness
Figure 4: Our method, KL-tracing using LRAS extracts better flow than other generative video models. (A) Deterministic models, such as CWM [bear_unifying_2023], often produce blurry predictions as they model a single, average future state. (B) Stable Video Diffusion lacks fine-grained controllability due to its coarse global latent code. Its clean and perturbed predictions differ in locations where the perturbation is not supposed to be carried to. (C) The Cosmos autoregressive world model lacks fine-grained controllability as it does not utilize pointers to denote the position of each token, making it challenging to prompt for flow extraction. (D) The LRAS model is highly controllable and has minimal differences between the clean and perturbed predictions. We use KL-tracing to compute the difference in logit instead of RGB space, obtaining sharp flow extractions.
@@
Figure 6: Large video models such as LRAS capture aspects of real-world dynamics that are challenging for specialized flow models relying on visual similarity or photometric loss. Specialized optical flow baselines, both supervised and self-supervised, struggle on complex, real-world scenes that are not fully captured by their training heuristics (Section 6). A deeper understanding of dynamics beyond feature/photometric similarity allows large video models such as LRAS to resolve motion in the presence of homogeneous objects, motion blur, and partial occlusion, among others.
6 Large video models capture aspects of real-world dynamics that are challenging for specialized flow models
Method: We compare KL-tracing with supervised flow methods such as RAFT [teed_raft_2020] and SEA-RAFT [wang_sea-raft_2024], as well as unsupervised methods such as Doduo [jiang_doduo_2023] and SMURF [stone_smurf_2021] (Table 3, Figure 6). We also compare our method to an alternative branch of approaches, multi-frame trackers, which typically use multiple frames of information to generate point tracks across time (Appendix C).
Findings: Powerful video models capture challenging aspects of real-world dynamics. For example, a globe undergoes in-place rotation with the query point on a textureless surface (Figure 1), or physical motion examples such as Newton’s cradle (Figure 6). These real-world dynamics are challenging for specialized flow models which are small and not trained on diverse, large-scale video datasets. Video models use generic training objectives, so they do not suffer from failure modes introduced by training heuristics used by specialized flow models. Supervised methods are trained on synthetic datasets as labeling real-world videos is expensive, and hence face a sim-to-real gap where they do not observe challenging real-world dynamics. Self-supervised methods such as Doduo [jiang_doduo_2023] and SMURF [stone_smurf_2021] use photometric loss for training, which results in challenges in predicting flow for frame pairs with significant differences in light intensity, and also enforce strong global consistency and smoothness constraints, which introduce failure modes for complex non-uniform dynamics that contain various magnitudes of flow across the same frame pair.
CWM [bear_unifying_2023] SVD [blattmann_stable_2023] Cosmos [nvidia_cosmos_2025] LRAS [lee_3d_2025]
@@
Stable Video Diffusion [blattmann_stable_2023] 74.7990
Cosmos (top 10% raster) (5MM, 2STD, $\times512$) [nvidia_cosmos_2025] 35.4338
Cosmos (overwrite 10% during rollout) (5MM, 2STD, $\times512$) [nvidia_cosmos_2025] 37.7552
Cosmos (provide full second frame) (5MM, 2STD, $\times512$) [nvidia_cosmos_2025] 66.5521
Table 2: KL-tracing with LRAS beats other video models. $M$ = multi-mask (average over multiple random masks), MS = multi-scale (zoom into image at multiple scales).
@@
AJ $\uparrow$ AD $\downarrow$ $\delta^{x}_{\text{avg}}\! \uparrow$ OA $\uparrow$ AJ $\uparrow$ AD $\downarrow$ $\delta^{x}_{\text{avg}}\! \uparrow$ OA $\uparrow$
@@
S RAFT [teed_raft_2020] 41.77 25.33 54.37 66.40 71.93 5.60 82.15 88.54
SEA-RAFT [wang_sea-raft_2024] 43.41 20.18 58.69 66.34 75.06 6.54 84.63 89.50
W Doduo [jiang_doduo_2023] 23.34 13.41 48.50 47.91 54.98 5.31 72.20 73.56
U SMURF [stone_smurf_2021] 30.64 27.28 44.18 59.15 65.81 6.81 80.57 87.91
DiffTrack [nam2025emergenttemporalcorrespondencesvideo] - - 46.90 - - - - -
CWM [bear_unifying_2023, venkatesh_counterfactual_2023] 15.00 23.53 26.30 76.63 28.77 11.64 41.63 84.93
@@
Table 3: TAP-Vid First: quantitative results on DAVIS and Kubric. Tracking starts when a point first appears and continues to the video end, thus involving large frame gaps. LRAS with KL-tracing outperforms two-frame baselines. “S” = supervised, “W” = weakly supervised, “U” = unsupervised.
@@ 7 Discussion
Taming generative video models with fine-grained controllability. A key finding of our work is that zero-shot prompting many mainstream generative video models for visual property extraction can be challenging due to their lack of fine-grained controllability. However, a series of design features can help tame these models: (1) predicting a distribution of future states rather than a deterministic average; (2) encoding each spatio-temporal patch independently; and (3) having a mechanism to decode individual parts of a frame in any given order. Currently, the LRAS [lee_3d_2025] framework was the only one we identified that contained all of these properties, but they could be adopted by other generative models as well, increasing the ease with which they can be visually prompted.
Tight coupling with the base generative model. The main advantage of using a generative model for flow is that it becomes easier to improve flow prediction quality on new domains. Any previously unrepresented domain can become “in-distribution” for KL-tracing, simply by extending the training data of the base generative model (e.g., LRAS). There is no need to obtain new, task-specific labels, which is particularly appealing for flow due to the intractable cost of human labeling for anything other than synthetic, computer-graphic type scenes. As modern video datasets continue to scale at a rapidly increasing rate, KL-tracing’s performance can improve in lockstep.
Distillation. Compared with real-time dense flow baselines [teed_raft_2020, wang_sea-raft_2024] or diffusion-based methods that build on feature correspondence [tang2023emergentcorrespondenceimagediffusion, nam2025emergenttemporalcorrespondencesvideo], KL-tracing can be slower due to the process of propagating interventions point-by-point. However, on challenging, real-world scenes, the quality of the extracted flow is notably better (Table 3). In fact, we can improve the Pareto frontier by taking the sparse labels that are extracted and distilling them into a faster architecture such as SEA-RAFT, thereby using KL-tracing as a much more scalable alternative to human labeling.
Future work. A natural next step is extending this test-time inference procedure to extract other visual intermediates such as object segments, material properties, or motion affordances. By moving away from fine-tuning our video models on labeled datasets to a paradigm of zero-shot prompting, the vision community can experience a shift akin to the one seen with large language models, moving from rigid representations constructed by fine-tuning toward dynamic, task-specific and adaptable ones extracted on the fly.
Overall, our results indicate that prompting controllable, self-supervised video models is a scalable and effective alternative to supervised or photometric-loss approaches for high-quality optical flow.
Acknowledgments and Disclosure of Funding
This work was supported by the following awards: Simons Foundation grant SFI-AN-NC-GB-Culmination-00002986-05, National Science Foundation CAREER grant 1844724, National Science Foundation Grant NCS-FR 2123963, National Science Foundation Grant RI 2211258, Office of Naval Research grant N00014-20-1-2589, ONR MURI N00014-21-1-2801, ONR MURI N00014-24-1-2748, and ONR MURI N00014-22-1-2740. We also thank the Stanford HAI, Stanford Data Sciences, the Marlowe team, and the Google TPU Research Cloud team for their computing support.
@@ Appendix A Additional details on methods
CWM [bear_unifying_2023] White Gaussian Frame 1, Partial Frame 2 RGB difference
SVD [blattmann_stable_2023] White Gaussian Noised latents of Frame 1 and 2 RGB difference
Cosmos [nvidia_cosmos_2025] White Gaussian Frame 1, Partial Frame 2 RGB difference
LRAS, RGB [lee_3d_2025] White Gaussian Frame 1, Partial Frame 2 RGB difference
@@ Appendix B Additional results
Figure 7: All evaluation settings for Cosmos [nvidia_cosmos_2025]) result in poor flow extractions. See Section 4.3 for a more detailed explanation of each result.
Appendix C Multi-frame trackers
We are primarily interested in evaluating a model’s ability to resolve motion from two frames, with methods and baselines chosen accordingly. While multi-frame trackers often achieve improved results due to aggregating and refining predictions globally, the two-frame setting directly evaluates a model’s core dynamics understanding, especially in more challenging real-world cases. We compare our method with supervised, multi-frame trackers run under the same two-frame constraint.
Method DAVIS
AJ $\uparrow$ $\delta^{x}_{\text{avg}}\! \uparrow$ OA $\uparrow$
TAP-Vid First
S CoTracker-v3 [karaev_cotracker_2024] 39.9 58.0 76.9
AllTracker [harley2025alltrackerefficientdensepoint] 54.9 66.0 78.0
U LRAS with KL-tracing (ours) 44.2 65.2 74.6
Table 5: TAP-Vid DAVIS results for two-frame trackers. Tracking starts when a point first appears and continues to the video end (large frame gaps). The supervised tracker baselines are natively multi-frame, but are evaluated in a two-frame setting for direct comparison.
The performance of multi-frame trackers is significantly impacted when removing context from intermediate frames. Notably, we find that AllTracker [harley2025alltrackerefficientdensepoint] is more robust, due to its approach of merging a two-frame, flow-type correspondence with a tracker-style global refinement across frames. Despite being completely training-free, our method shows comparable performance with the latest trackers in this challenging setting (Table 5).